If A =

(i 0
0 i)
, where i =
√−1, find A^2
and A^4
(Matrices)
1
Expert's answer
2020-05-27T17:13:13-0400
"A=\begin{pmatrix}
 i & 0\\
 0 & i
\end{pmatrix}"

"A^2=\begin{pmatrix}
 i & 0\\
 0 & i
\end{pmatrix}\begin{pmatrix}
 i & 0\\
 0 & i
\end{pmatrix}="

"=\begin{pmatrix}
 i^2+0 & 0+0\\
 0+0 & 0+i^2
\end{pmatrix}=\begin{pmatrix}
 -1 & 0\\
 0 & -1
\end{pmatrix}"

"A^4=\begin{pmatrix}
 -1 & 0\\
 0 & -1
\end{pmatrix}\begin{pmatrix}
 -1 & 0\\
 0 & -1
\end{pmatrix}="

"=\begin{pmatrix}
 (-1)^2+0 & 0+0\\
 0+0 & 0+(-1)^2
\end{pmatrix}=\begin{pmatrix}
 1 & 0\\
 0 & 1
\end{pmatrix}"